1. given the function H(z)=2x+6 find eac of the following
h(8)=
H(-5)=
H (6.6)=
2.Write the equation of a line containing the given point and parallel to the given line (-9,2); 7x=3y+5
3. find an equation of a the line having the given slope and containing the given point m=-8, (7,8)

Answer by stanbon(75887)

1. given the function H(x)=2x+6 find each of the following
h(8)= 2*8+6 = 22
H(-5)= 2*-5+6 = -4
H (6.6)= 2*6.6+6 = 19.2
-----------------------------------
2.Write the equation of a line containing the given point and parallel to the given line (-9,2); 7x=3y+5
The line has slope m=7/3
Form: y=mx+b ; you have x,y, and m; solve for b
2 = (7/3)(-9) +b
2 = -21 + b
b = 23
EQUATION:
y = (7/3)x+23
=========================
3. find an equation of a the line having the given slope and containing the given point m=-8, (7,8)
8 = -8*7 + b
b = 8+56 = 64
EQUATION:
y = -8x+64
==================
Cheers,
Stan H.